OpenStudy (anonymous):

Evaluate the limit lim (25-b)/(5-√b) b-->25

OpenStudy (turingtest):

factor the numerator; something will cancel

OpenStudy (turingtest):

remember the difference of squares formula\[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

i tried to factor but it didn't work?

OpenStudy (turingtest):

what did you factor the numerator into?

OpenStudy (anonymous):

5.5-b?

OpenStudy (turingtest):

not sure where you got 5.5 from use the formula for difference of squares I worte above\[a^2-b^2=(a-b)(a+b)\]what is \(a\) in your case? what is \(b\) ?

OpenStudy (turingtest):

wrote

OpenStudy (anonymous):

a=5 b=b?

OpenStudy (turingtest):

a=5 yes but in your formula b^2=b (bad notation choice on my part, sorry, different b's) so when you factor it a=5 and b=√b

OpenStudy (anonymous):

(5-b)(5+b)

OpenStudy (turingtest):

let me rephrase the difference of squares formula to avoid confusion...

OpenStudy (turingtest):

in the form \[p^2-q^2\]we can factor this into\[(p-q)(p+q)\]you have, in the numerator\[25-b\]so what is p^2 ? what is q^2 ? based on that, what is p and what is q ? then write your expression

OpenStudy (anonymous):

is it (5-b)(5+b)?

OpenStudy (turingtest):

not quite what is q^2 in your expression?

OpenStudy (anonymous):

5?

OpenStudy (turingtest):

no, that is p let's compare your equation with the form for difference of squares\[p^2-q^2=(p-q)(p+q)\]\[25-b=?????\]so we can see that \(25=p^2\), which means that \(p=\pm25\) now what about \(q^2\) what does that correspond to in your formula?

OpenStudy (turingtest):

typo above *which means that \(p=\pm5\)

OpenStudy (anonymous):

ok how about q^2?

OpenStudy (anonymous):

b^2

OpenStudy (turingtest):

let us make a comparison